twice the sum of 4 and three times some number is 34. whats the number?
Answer by unlockmath(1122)

Hello,
Let's put this into the following. (Hint:In word problems the word "is" usually means equal.)
2(4+3x)=34 Multiply this out will give us:
8+6x=34 Subtract 8 from both sides will be:
6x=26 Divide 6 into both sides will be:
x = 26/6 or 4 1/3
RJ Toftness
www.math-unlock.com